# Towards understanding evolution of science through language model series

### v1

@@ V-B1 Experimental setup
The arXiv citation network is actually a temporal (growing) network, as we know the publication year of each paper. This raises the question of how different models perform in a more challenging setting of prospective link prediction, where we train a classifier using links and non-links observed during an earlier period and then apply it to predict the appearance of future links. Formally, let $G_{t_{0}}=(V_{t_{0}},E_{t_{0}})$ denote the citation network observed at year $t_{0}$, where the node set $V_{t_{0}}=\{v|t_{v}\leq t_{0}\}$ represents all the arXiv papers published until $t_{0}$ and $E_{t_{0}}=\{(u,v)|u,v\in V_{t_{0}}\}$ is the set of citation links between them. With $G_{t_{0}}$, we use GraphSAGE and MLP to model interactions between embeddings of two arXiv papers in $V_{t_{0}}$, following the same process as in the static network case. We then apply the resulting architecture to predict links appeared in $t_{0}+1$. Specifically, we denote the test dataset as $\mathcal{D}_{\text{test}}=(B_{t_{0}+1}^{+},B_{t_{0}+1}^{-})$, where $B_{t_{0}+1}^{+}$ contains all the observed citations (positive examples) from papers published in $t_{0}+1$ to papers published until $t_{0}+1$, i.e., $B_{t_{0}+1}^{+}=\{(u,v)|t_{u}=t_{0}+1,t_{v}\leq t_{0}+1\}$, and $B_{t_{0}+1}^{-}$ is the set of non-existent links (negative examples) that are generated by, for each observed link $(u,v)\in B_{t_{0}+1}^{+}$, randomly choosing a different endpoint $v^{\prime}$ published until $t_{0}$, i.e., $B_{t_{0}+1}^{-}=\{(u,v^{\prime})|t_{u}=t_{0}+1,t_{v}^{\prime}\leq t_{0},(u,v)\not\in B_{t_{0}+1}^{+}\}$.
[…]
V-B2 Experimental results
TABLE IV: Experimental results for link prediction in the temporal arXiv citation network.
Model AUC-ROC Accuracy Precision Recall F1 Score
GraphSAGE + Random 0.660 0.582 0.651 0.352 0.457
GraphSAGE + Major category 0.756 0.576 0.809 0.198 0.318
GraphSAGE + Subcategory 0.965 0.900 0.942 0.853 0.896
GraphSAGE + BERT 0.951 0.846 0.931 0.747 0.829
GraphSAGE + BioBERT 0.960 0.866 0.931 0.790 0.855
GraphSAGE + SciBERT 0.972 0.880 0.956 0.796 0.869
GraphSAGE + BioMedBERT 0.962 0.871 0.940 0.792 0.860
GraphSAGE + AnnualBERT 0.988 0.939 0.968 0.907 0.937
@@ V-B2 Experimental results
Table IV presents the experimental results, comparing the performances of the 8 different methods for generating the $H^{(0)}$ representations. Firstly, the two methods that provide insufficient information (i.e., random and major category) perform poorly across all the metrics on the temporal network, as there lacks topological information that could be leveraged during the prediction process. By contrast, the subcategory-based method yields surprisingly strong results, even surpassing BERT and BioBERT in terms of AUC-ROC. This impressive performance may be attributed to the specificity of the subcategories, with a total of 172 distinct subcategories providing highly targeted information. Our AnnualBERT model still stands out distinctly when comparing with the other BERT-based models, achieving the highest scores in AUC-ROC (0.988), accuracy (0.939), recall (0.907), and F1 Score (0.937), underscoring the effectiveness of domain-specific pretraining in capturing richer semantic representations.
VI Mining AnnualBERT
@@
In the previous two sections, we have demonstrated the validity of our AnnualBERT series of models ${\mathcal{M}_{t}}$ and their superior performances than the other scientific BERT models for tasks specific to the arXiv domain. In this section, we are interested in mining the weights of these models, as we reason that each of the models serves as a condensed representation of the scientific literature over a specific period, the mining of which would enable us to understand the changing landscape of science as captured in the arXiv corpus.
Figure 4: PCA visualizations of model weights in different layers.
To motivate, we examine if and how the model weights change when it keeps adapting to new corpora published over time. To this end, we select various layers to visualize their weights. For instance, in the word embedding layer, we transform each model’s $53100\times 768$ matrix associated with that layer into a one-dimensional vector, resulting in a $13\times 40780800$ matrix. Subsequently, we employ principal component analysis (PCA) to reduce the dimensionality of the flattened word embedding weights to two dimensions. The same procedure is applied to other layers to obtain visualizations shown in Fig. 4, from which we draw two observations. Firstly, for the majority of layers, two components account for over 60% of the variance, while in the case of the word embedding layer, they explain almost 90% of the variance, indicating that the model weights are highly redundant and possess distillability [57]. Secondly, Fig. 4 also shows that, across layers, the model’s weight space appears to have a distinct pattern, with all the weight points lying on a quadratic curve, implying that the model parameters exhibit observable changes when the training corpora are organized by year. These results support the hypothesis that continual pretraining on corpora formed by time leads to constant adaptations in the model’s knowledge representation. This finding is important for understanding how changes in scientific literature are encapsulated within the evolving parameters of domain-specific language models like AnnualBERT.
VI-A Temporal adaption
VI-A1 Experimental setup
Figure 5: Summary F1 matrices $\widehat{P}_{t,\tau}$ showing how AnnualBERT models $\mathcal{M}_{t}$ perform in prediction tasks on papers from different data years $\tau$. (a–b) Major category prediction based on (a) entire abstracts and (b) only second half of abstracts. (c) Subcategory classification; (d) Cross-field identification.
@@ VI-A1 Experimental setup
VI-A2 Experimental results
@@ VI-B Interpolation of models
$\mathcal{M}_{t_{0}}^{\Delta t}=\frac{1}{2}\left(\mathcal{M}_{t_{0}-\Delta t}\oplus\mathcal{M}_{t_{0}+\Delta t}\right)\,.$ (5)
@@
Focusing on $t_{0}=2014$, we have 6 interpolated models by varying $\Delta t$, and Fig. 6 presents the average F1 scores of these models together with $\mathcal{M}_{2014}$ for the 3 tasks. We also display the $p$-values of the Mann-Whitney U tests comparing F1 scores of each interpolated model with $\mathcal{M}_{2014}$. Firstly, we observe that across the 3 tasks, all the interpolated models perform well and exhibit comparable performances with the real model. To further demonstrate that both $\mathcal{M}_{t_{0}-\Delta t}$ and $\mathcal{M}_{t_{0}+\Delta t}$ are necessary in creating the interpolated model, we note that using either of them would not generate similar performance as $\mathcal{M}_{t_{0}}$ (see Figs. 5a, c, d). Furthermore, we conduct additional experiments where we replace $\mathcal{M}_{2008}$ with a purely random model $\mathcal{M}_{2008}^{\text{random}}$ that has the same mean and variance as $\mathcal{M}_{2008}$, and interpolating between $\mathcal{M}_{2020}$ and $\mathcal{M}_{2008}^{\text{random}}$ yields significantly worse performances than $\mathcal{M}_{2014}$ for the 3 tasks (F1 scores are 0.66, 0.07, and 0.79 respectively, with all $p$-values below 0.05.).
@@
The second observation from Fig. 6 is that as $\Delta t$ increases (interpolating between two models farther way temporarily), the performance gap between $\mathcal{M}_{2014}^{\Delta}$ and $\mathcal{M}_{2014}$ tend to increase, although lack of statistical significance. This may imply that the model weights encode temporal information via continual training, and it is feasible to edit the model to fulfil the desired task within a specific time frame.
@@
Figure 6: Average F1 scores of $\mathcal{M}_{t_{0}}$ and interpolated models (denoted as $\text{Mix}_{t_{1},t_{2}}$) between $\mathcal{M}_{t_{1}}$ and $\mathcal{M}_{t_{2}}$. (a) Major category classification; (b) Subcategory classification; (c) Cross-field identification.
VI-C Random shuffle experiment
Finally, having examined temporal adaptability and interpolation of AnnualBERT, we explore the root cause of these behaviors—training corpus—and study how it affects the learning and forgetting behaviors by designing experiments on synthetic datasets. In particular, we first create a new, shuffled corpus, where we randomly shuffle the sequence of the words in a sentence from the abstracts in 2020. We then rerun continual training of $\mathcal{M}_{2020}$ on this shuffled corpus for 5 epochs, resulting in a new model $\mathcal{M}_{2020}^{(\text{s})}$. For comparison, we also repeat the same procedure but on the original, unaltered abstracts to obtain $\mathcal{M}_{2020}^{(\text{o})}$. In the shuffled case, we force the model to adapt to a semantically unmeaningful corpus and consequently to forget the learned semantics, whereas in the unaltered case, the model reinforces the learning process. Next, we evaluate the extent of learning and forgetting in the same manner as in § VI-A, by sampling from shuffled abstracts a training set and a test set, denoted as $\mathcal{D}^{(\text{s})}$, and similarly from unshuffled abstracts $\mathcal{D}^{(\text{o})}$. We then test the classification performances of the two models on the two datasets and get the following accuracy matrix:
$\mathcal{D}^{(\text{o})}$ $\mathcal{D}^{(\text{s})}$
$\mathcal{M}_{2020}^{(\text{o})}$ $A_{o,o}$ $A_{o,s}$
$\mathcal{M}_{2020}^{(\text{s})}$ $A_{s,o}$ $A_{s,s}$
For dataset $\mathcal{D}^{(\text{o})}$, we quantify performance deterioration by calculating relative accuracy to the baseline:
$\widetilde{A}^{\text{o}}=\frac{A_{s,o}-A_{o,o}}{\text{baseline}}\,.$ (6)
Here baseline is the accuracy by random guessing the label:
$\text{baseline}=\sum_{i}\left(\frac{N_{i}}{N}\right)^{2}\,,$ (7)
where $N_{i}$ is the number of samples of class $i$ and $N$ is the total samples. For dataset $\mathcal{D}^{(\text{s})}$, the relative accuracy $\widetilde{A}^{\text{s}}$ is defined similarly.
Table V presents $\widetilde{A}^{\text{o}}$ and $\widetilde{A}^{\text{s}}$ for the 3 tasks. It indicates a persistent performance shift across tasks: The model $\mathcal{M}_{2020}^{(\text{s})}$ trained on a shuffled corpus under-performs on real data but over-performs on shuffled data, showcasing the model’s adaptation to the disordered text and exhibition of semantic forgetting. This phenomenon is most pronounced for subcategory prediction, where the model critically relies on the semantically meaningful text to learn the subtleties useful in identifying fine grained categories, the destroying of which causes the model under-perform severely. This result is also consistent with Fig. 5c showing that the model needs sustained training instances to have a reliable representation of a particular subcategory.
TABLE V: Relative accuracy of 3 prediction tasks on the original and shuffled abstracts.
Task $\widetilde{A}^{\text{o}}$ $\widetilde{A}^{\text{s}}$
Major category prediction -0.002 0.030
Subcategory prediction -0.779 0.392
Cross-field identification -0.014 0.007
VII Conclusion
@@
In this work, we presented a new method for pretraining a series of models by organizing the corpora chronologically and introduced AnnualBERT, a series of language models for academic text with different years. This series of language models allows for the tracking of scientific documents across various time periods, facilitating the representation of knowledge over time. We validated our models by fine-tuning them on several standard benchmark datasets, demonstrating that our method achieved a comparable performance to similar models. Moreover, we employed a domain-specific link prediction task on the arXiv dataset to evaluate the performance of our model, and the results confirmed that our model achieved the state of the art performance in the domain-specific context. To comprehend how models adjust to the evolution of scientific discourse over extended periods, we visualized the model weights and employed a probing task classifying the arXiv corpus to quantify the model’s behavior in terms of learning and forgetting as time progresses. Our findings revealed that our model not only exhibits performance on par with the other domain-specific BERT models in standard natural language processing tasks but also captures and reflects the unique linguistic characteristics of the corresponding publication year.
Acknowledgment

### v2

@@ 5.2.1. Experimental setup
The arXiv citation network is actually a temporal (growing) network, as we know the publication year of each paper. This raises the question of how different models perform in a more challenging setting of prospective link prediction, where we train a classifier using links and non-links observed during an earlier period and then apply it to predict the appearance of future links. Formally, let $G_{t_{0}}=(V_{t_{0}},E_{t_{0}})$ denote the citation network observed at year $t_{0}$, where the node set $V_{t_{0}}=\{v|t_{v}\leq t_{0}\}$ represents all the arXiv papers published until $t_{0}$ and $E_{t_{0}}=\{(u,v)|u,v\in V_{t_{0}}\}$ is the set of citation links between them. With $G_{t_{0}}$, we use GraphSAGE and MLP to model interactions between embeddings of two arXiv papers in $V_{t_{0}}$, following the same process as in the static network case. We then apply the resulting architecture to predict links appeared in year $t_{0}+\Delta t$, where $\Delta t$ denotes the number of years into the future. Specifically, we denote the test dataset as $\mathcal{D}_{\text{test}}=(B_{t_{0}+\Delta t}^{+},B_{t_{0}+\Delta t}^{-})$, where $B_{t_{0}+\Delta t}^{+}$ contains all the observed citations (positive examples) from papers published in year $t_{0}+\Delta t$ to papers published until $t_{0}+\Delta t$, i.e., $B_{t_{0}+\Delta t}^{+}=\{(u,v)|t_{u}=t_{0}+\Delta t,t_{v}<t_{0}+\Delta t\}$, and $B_{t_{0}+\Delta t}^{-}$ is the set of non-existent links (negative examples) that are generated by, for each observed link $(u,v)\in B_{t_{0}+\Delta t}^{+}$, randomly choosing a different endpoint $v^{\prime}$ published until $t_{0}$, i.e., $B_{t_{0}+\Delta t}^{-}=\{(u,v^{\prime})|t_{u}=t_{0}+\Delta t,t_{v^{\prime}}<t_{0},(u,v^{\prime})\not\in B_{t_{0}+\Delta t}^{+}\}$.
In our experiment, we consider $t_{0}=2014$ and vary $\Delta t$ from 1 to 6, thus predicting future links appeared from 2015 to 2020. The network $G_{t_{0}}$ has $\left|V_{t_{0}}\right|=637,740$ nodes and $\left|E_{t_{0}}\right|=3,999,781$ links. We evaluate the performances of the above mentioned 8 ways of obtaining initial node embeddings $H^{(0)}$ by testing on the $\mathcal{D}_{\text{test}}$ datasets. For our method, we take AnnualBERT $\mathcal{M}_{t_{0}}$ for feature encoding. Here we do not test the one-time trained model $\overline{\mathcal{M}}$, since it contains future information.
5.2.2. Experimental results
@@
Fig. 4 presents the experimental results, comparing the performances of the 8 different methods for generating the $H^{(0)}$ representations. Overall, the performance of most methods declines as the time shift $\Delta t$ increases, aligning with the expectation of a growing challenge of predicting future links over a longer period. However, the performance decay is largely driven by the deterioration of recall (Fig. 4D), while the precision remains stable (Fig. 4C). This highlights the difficulty in retrieving pairs of papers with citation relationships (true positives), potentially pointing to the emergence of scientific innovation resulted from making novel combinations (Uzzi et al., 2013; Ke, 2020). This could serve as a future direction to explore the relationship between predictability of citation pairs and scientific impact. Looking at different methods, the two encoders with limited information (i.e., random and major category) perform poorly across all metrics, as they lack the contextual and structural information necessary for effective link prediction. By contrast, the subcategory-based method yields surprisingly strong results, even surpassing most methods except AnnualBERT in terms of recall and F1. This impressive performance may be attributed to the specificity of the subcategories, with a total of 172 distinct subcategories providing highly targeted information. Our AnnualBERT model still consistently outperforms with a significantly large gap when comparing with the other BERT-based models, achieving the highest scores in AUC-ROC, accuracy, recall, and F1 Score, underscoring the effectiveness of domain-specific pretraining in capturing richer semantic and temporal representations, making it particularly robust for long-term link prediction.
Figure 4. Experimental results for link prediction in the temporal arXiv citation network.
6. Mining AnnualBERT
@@
In the previous two sections, we have demonstrated the validity of our AnnualBERT series of models ${\mathcal{M}_{t}}$ and their superior performances than the other scientific BERT models for tasks specific to the arXiv domain. In this section, we are interested in mining these models, as we reason that each of the models serves as a condensed representation of the scientific literature over a specific period, the mining of which would enable us to understand the changing landscape of science as captured in the arXiv corpus.
6.1. Temporal adaption
6.1.1. Motivation
As mentioned above, our choice of using whole word rather than subword as token is motivated by the fact that words are the most basic constituent unit for ideas, and the probabilistic feature of BERT to predict the masked token given its context (other tokens in the sentence) allows us to analyze the dynamics of research by examining how the probability of each word change over time, as assessed by our AnnualBERT model series. For example, consider the sentence shown in Fig. 5 about edge representation in image models, we mask “ResNet” and ask each of the AnnualBERT models $\mathcal{M}_{t}$ to output the most likely tokens and their associated probabilities (Fig. 5A). For comparisons, we also test with the one-time trained $\overline{\mathcal{M}}$ (Fig. 5B) and SciBERT (Fig. 5C). For SciBERT, as it uses subword tokenization, meaningful ideas like “ResNet”, “AlexNet”, “GoogLeNet”, and “VGG” are all broken into subwords like “alex”, “google”, and “vg”. Consequently, SciBERT is not able to make the correct guess, leaving “cnn” as the most likely token and “sub-ideas” like “alex” and “google” small probabilities. By contrast, $\overline{\mathcal{M}}$ adopts whole-word tokenization, and most of the top choices are image models. Yet, the probability of “ResNet” (0.45) does not overwhelmingly dominate that of “AlexNet” (0.18), indicating that the latter one is still a viable choice. Finally, annually trained models $\mathcal{M}_{t}$ consider further the temporal dimension and reveal the dynamics of the competition between AlexNet and ResNet. In summary, this example highlights the rationale and the advantage of whole-word tokenization for the analysis of scientific texts where new terminologies are constantly emerging.
Figure 5. The probabilities of top tokens given by different models, for the sentence shown at the top.
This analysis raises a question: How do the model’s internal weights evolve to capture such contextual shifts in word usage over time? To explore this, we examine how the model weights change as AnnualBERT adapts to new corpora published annually. Specifically, we focus on visualizing the weights of different layers in the model. To this end, we select various layers to visualize their weights. For instance, in the word embedding layer, we transform each model’s $53100\times 768$ matrix associated with that layer into a one-dimensional vector, resulting in a $13\times 40780800$ matrix. Subsequently, we employ principal component analysis (PCA) to reduce the dimensionality of the flattened word embedding weights to two dimensions. The same procedure is applied to other layers to obtain visualizations shown in Fig. 6, from which we draw two observations.
  1. (1)
    Redundancy in model weights: For most layers, the first two principal components explain over 60% of the variance, with the word embedding layer exhibiting even higher redundancy (nearly 90%). This suggests that the model weights are highly compressible, aligning with findings on the distillability of large language models (Liu et al., 2022a).
  2. (2)
    Temporal trajectory of weights: The weight space across layers exhibits a distinct quadratic pattern, with model weights from different years lying along a smooth curve. This indicates that the model parameters undergo systematic changes as they adapt to new corpora, implying the temporal evolution of scientific language.
These results support the hypothesis that continual pretraining on corpora formed by time leads to constant adaptations in the model’s knowledge representation. This finding is important for understanding how changes in scientific literature are encapsulated within the evolving parameters of domain-specific language models like AnnualBERT.
Figure 6. PCA visualizations of model weights in different layers.
6.1.2. Experimental setup
@@
6.1.3. Experimental results
Figure 7. Summary F1 matrices $\widehat{P}_{t,\tau}$ showing how AnnualBERT models $\mathcal{M}_{t}$ perform in prediction tasks on papers from different data years $\tau$. (a–b) Major category prediction based on (a) entire abstracts and (b) only second half of abstracts. (c) Subcategory classification; (d) Cross-field identification.
@@ 6.2. Interpolation of models
(5) $\mathcal{M}_{t_{0}}^{\Delta t}=\frac{1}{2}\left(\mathcal{M}_{t_{0}-\Delta t}\oplus\mathcal{M}_{t_{0}+\Delta t}\right)\,.$
@@
Focusing on $t_{0}=2014$, we have 6 interpolated models by varying $\Delta t$, and Fig. 8 presents the average F1 scores of these models together with $\mathcal{M}_{2014}$ for the 3 tasks. We also display the $p$-values of the Mann-Whitney U tests comparing F1 scores of each interpolated model with $\mathcal{M}_{2014}$. Firstly, we observe that across the 3 tasks, all the interpolated models perform well and exhibit comparable performances with the real model. To further demonstrate that both $\mathcal{M}_{t_{0}-\Delta t}$ and $\mathcal{M}_{t_{0}+\Delta t}$ are necessary in creating the interpolated model, we stress that using either of them would not generate similar performance as $\mathcal{M}_{t_{0}}$ (see Figs. 7a, c, d). Furthermore, we conduct additional experiments where we replace $\mathcal{M}_{2008}$ with a purely random model $\mathcal{M}_{2008}^{\text{random}}$ that has the same mean and variance as $\mathcal{M}_{2008}$, and interpolating between $\mathcal{M}_{2020}$ and $\mathcal{M}_{2008}^{\text{random}}$ yields significantly worse performances than $\mathcal{M}_{2014}$ for the 3 tasks (F1 scores are 0.66, 0.07, and 0.79 respectively, with all $p$-values below 0.05.).
@@
The second observation from Fig. 8 is that as $\Delta t$ increases (interpolating between two models farther way temporarily), the performance gap between $\mathcal{M}_{2014}^{\Delta}$ and $\mathcal{M}_{2014}$ tend to increase, although lack of statistical significance. This may imply that the model weights encode temporal information via continual training, and it is feasible to edit the model to fulfill the desired task within a specific time frame.
@@
Figure 8. Average F1 scores of $\mathcal{M}_{t_{0}}$ and interpolated models (denoted as $\text{Mix}_{t_{1},t_{2}}$) between $\mathcal{M}_{t_{1}}$ and $\mathcal{M}_{t_{2}}$. (a) Major category classification; (b) Subcategory classification; (c) Cross-field identification.
7. Conclusion
@@
In this work, we presented a new method for pretraining a series of models by organizing the corpora chronologically and introduced AnnualBERT, a series of language models for academic text with different years. This series of language models allows for the tracking of scientific documents across various time periods, facilitating the representation of knowledge over time. We validated our models by fine-tuning them on several standard benchmark datasets, demonstrating that our method achieved a comparable performance to similar models. Moreover, we employed domain-specific link prediction tasks on the arXiv dataset to evaluate the performance of our model, and the results confirmed that our model achieved the state of the art performance in the domain-specific context. To comprehend how models adjust to the evolution of scientific discourse over extended periods, we visualized the model weights and employed a probing task classifying the arXiv corpus to quantify the model’s behavior in terms of learning and forgetting as time progresses. Our findings revealed that our model not only exhibits performance on par with the other domain-specific BERT models in standard natural language processing tasks but also captures and reflects the unique linguistic characteristics of the corresponding publication year.
Acknowledgements.